MOLECULAR ROTATION IN SUPERFLUID ^{4}He NANODROPLETS: THE VALIDITY OF A HYDRODY-NAMIC MODEL

\underline{CARLO CALLEGARI}, ANDRÉ CONJUSTEAU, IRENE REINHARD, KEVIN K. LEHMANN, and GIACINTO SCOLES, Department of Chemistry, Princeton University, Princeton NJ 08544; FRANCO DALFOVO, Dipartimento di Matematica e Fisica, Università Cattolica del Sacro Cuore, I-25121 Brescia (Italy).

In order to accurately predict the effective moments of inertia (I_{eff}) of linear molecules rotating in superfluid ^{4}He droplets, we have recently devised a superfluid hydrodynamic model. In the present implementation, the helium density profile induced by the He-molecule interaction potential is first calculated at the Density Functional level, and then used as the input of the hydrodynamic equation for the irrotational motion of a viscousless fluid. The kinetic energy of the fluid (E_{k}) is then used to calculate I_{eff} via: E_{k} = \frac{1}{2} ΔI ω^{2}, where ω is the angular velocity of the molecule, and ΔI is the difference between I_{eff} and the moment of inertia of the bare molecule.^{a} The model relies on the assumption that the density of the fluid in the rotating frame of reference is independent of ω and can therefore be calculated in the limit of a static molecule (adiabatic following approximation). The validity of this approximation, and its first-order corrections will be discussed. The advantages and limitations of Density Functional Theory for calculating the true helium density will also be addressed, in a comparison with Quantum Monte Carlo results which recently became available for HCN and its oligomers.^{b}

aC. Callegari, A. Conjusteau, I. Reinhard, K.K. Lehmann, G. Scoles, and F. Dalfovo, Phys. Rev. Lett. 83, 5058 (1999); ibid.

84, 1848(E) (2000).

bE. Draeger and D. M. Ceperley, private communication.